# Simple-RF: Regularizing Sparse Input Radiance Fields with Simpler Solutions

### v2

@@ C.2. Visualization of Depth Reliability Masks
In Fig. 27, we present visualizations that motivate the design of our augmentations in Simple-NeRF, namely the smoothing and Lambertian augmentations. We train our model without augmentations and the individual augmentations separately with only $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{sd}}$ for 100k iterations. Using the depth maps predicted by the models for an input training view, we determine the mask that indicates which depth estimates are more accurate, as explained in Sec 4.4 of the main paper Verify if this is correct. For two scenes from the LLFF dataset, we show an input training view and focus on a small region to visualize the corresponding masks.

### v4

@@ A.1. Masked Evaluation
A.2. Difference in Evaluation Scores
We note that the evaluation scores of the comparison models differ from those reported in their respective papers due to the following reasons.
  • As noted above, we mask out the regions not visible in at least two of the input views during evaluation.
  • The number of train and test views used for experiments vary across the different papers. While they’re uniformly picked in certain cases, they are hard-coded in other cases. Further, different papers use different resolution for the input images. For a reliable comparison across different models, we fix the same train and test views as well as the resolution and re-run all the benchmarked models using the author-provided source code and report the results.
  • Finally, the SSIM scores vary across different implementations (Venkataramanan et al., 2021). Similarly, the LPIPS scores vary depending on the backbone used (Alex-net vs VGG-net). We use scikit-image implementation of SSIM and Alex-net backbone for LPIPS.
@@ C.2. Visualization of Depth Reliability Masks
In Fig. 29, we present visualizations that motivate the design of our augmentations in Simple-NeRF, namely the smoothing and Lambertian augmentations. We train our model without augmentations and the individual augmentations separately with only ${\mathcal{L}_{ph}}$ and ${\mathcal{L}_{sd}}$ for 100k iterations. Using the depth maps predicted by the models for an input training view, we determine the mask that indicates which depth estimates are more accurate, as explained in Sec 4.1.4 of the main paper. For two scenes from the LLFF dataset, we show an input training view and focus on a small region to visualize the corresponding masks.